Determine the root-mean-square speed of CO2 molecules that have an average kinetic energy of
4.21 × 10–21 J per molecule.
1
Expert's answer
2020-06-05T05:12:09-0400

The root mean square speed formula is

"E=\dfrac{1}{2}mv^2"

1 mole of CO2 =44.0 g or 0.044 kg

If energy is in form of Joules then, mass should be in kilograms.

"average mass of CO2 =\dfrac{mass ofCO2}{Avagadros constant}"


"=\dfrac{4.40*10^{-2}kg}{6.02*10^{23}kg}" = 7.31* 10^-26 kg

since"E=\dfrac {1}{2}mv^2 then,"

v^2=2 E/m


="2\dfrac{4.2*10^{-21}J}{7.31*10^{-26}kg}=115184.68"


V ="\sqrt{115184.68}"

= 339

therefore,root-mean-square speed is 339 m/s